Solve (D2+2D+2)y=0 Initial Conditions: y(0)=1, y'(0)=2

Solve (D2+2D+2)y=0

Initial Conditions: y(0)=1, y'(0)=2
Answer #1

Griven differential equatio CD?+20+2)y=o no 1. The equation is in the form of frowy =Rittis]. which is higher order different(Here a = b = 2, C=2] 0 -2 √(2) ² 4(1912) 211) Þ = -2 /4 -8 2 = -2174 o= - 212i D = Fifi) D = -lt i(LL.B) - dars and pri De tui chere? LE-1 &fio) L pai) ... The complemenefary function = Cof = axe, cospet Chings) = 184CC, COD+20+2 a. The general soluation is 962) = Yetyp 402) Ecet p. I %(= ēkie, consect Sfinx) +0 402) = 768, cosx ts finely Given. Sin(o)= 0 icoa = 1 é ºri subftiterte c, value in 1. equation @ we get You = ēx (1) cosxt sting) (93) = 2^(Ce/x + , in 2-9Here e- êt and ane w=cosset&rina de) = zz de les et spinat corzaçmeyecek) y (x) = ēxd cosxf Srina) + Colts dx 489 = 7 X asiyco) = 7 °C, cosco) -sin())-ě °C.1o+c, finco) 2 -1 (2011-0)-1(1+(0)) 2 =@2-0). I to) 2= ez-ici) 2 = CR- 2+1= CZ 2 = 37 subs